OpenStudy (anonymous):

Will give medal. What are the domain and range of the following function: f(x)=x^2-2x+2

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

it is a polynomial, so domain is all real numbers

OpenStudy (anonymous):

because the graph is a parabola that opens up, the range will be from the second coordinate of the vertex to infinity, so you need to find the vertex

OpenStudy (anonymous):

the first coordinate of the vertex of \(y=ax^2+bx+c\) is \(-\frac{b}{2a}\) you have \(f(x)=x^2-2x+2\) so the first coordinate of the vertex is \(-\frac{-2}{2\times 1}=1\)

OpenStudy (anonymous):

the second coordinate of the vertex is \(f(1)\) compute that number

OpenStudy (anonymous):

will it be domain=all real numbers, and range equal y>=1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok. Thank you. I've been trying to figure that out for awhile now. :)

OpenStudy (anonymous):

yw